# ControlNet+++++ +: Improving Conditional Controls with Efficient Consistency Feedback Project Page: liming-ai.github.io/ControlNet_Plus_Plus

### v1

@@ 5 Conclusion
In this paper, we demonstrate from both quantitative and qualitative perspectives that existing works focusing on controllable generation still fail to achieve precise conditional control, leading to inconsistency between generated images and input conditions. To address this issue, we introduce ControlNet++, it explicitly optimizes the consistency between input conditions and generated images using a pre-trained discriminative reward model in a cycle consistency manner, which is different from existing methods that implicitly achieve controllability through latent diffusion denoising. We also propose a novel and efficient reward strategy that calculates consistency loss by adding noise to input images followed by single-step denoising, thus avoiding the significant computational and memory costs associated with sampling from random Gaussian noise. Experimental results under multiple conditional controls show that ControlNet++ significantly improves controllability without compromising image quality, offering new insights into the field of controllable generation.
6 More Visualization
@@ 7 Overview of Supplementary
  • Section 10: Effectiveness of conditioning scale of existing methods.
  • Section 11: Broader impact and limitation.
@@ 8.1 Dataset Details
Considering that the training data for ControlNet [63] has not been publicly released, we need to construct our training dataset. In this paper, we adhere to the dataset construction principles of ControlNet [63], which endeavor to select datasets with more accurate conditional conditions wherever possible. Specifically, for the segmentation condition, previous works have provided datasets with accurately labeled segmentation masks [67, 68, 4]. Therefore, we opt to train our model using these accurately labeled datasets following ControlNet [63]. For the Hed, LineArt edge tasks, it is challenging to find datasets with real and accurate annotations. As a result, following ControlNet [63], we train the model using the MultiGen20M dataset [37], which is annotated by models, to address this issue. Regarding the depth task, existing datasets include masks of certain pixels as having unknown depth values, making them incompatible with the current ControlNet pipeline. Therefore, we also adapt the MultiGen20M depth dataset, which is similar to the dataset constructed by ControlNet [63]. In terms of the canny edge task, no human labels are required in the process, so we also adapt the MultiGen20M dataset. We provide details of the datasets in Tab. 5.
8.2 Reward Model and Evaluation Details
In general, we deliberately choose slightly weaker models as the reward model and opt for stronger models for evaluation. This practice not only ensures the fairness of the evaluation but also helps to determine whether performance improvements result from alignment with the reward model’s preferences or from a genuine enhancement in controllability. While such an approach is feasible for some tasks (Segmentation, Depth), it becomes challenging to implement for others (Hed, Canny, LineArt Edge) due to the difficulty in finding two distinct reward models. In such cases, we use the same model as both the reward model and the evaluation model. We utilize standard evaluation schemes from their respective research fields to evaluate the input conditions and extracted conditions from the generated images, as demonstrated in Sec. 4.1. We use the same Hed edge detection model and LineArt edge detection model as ControlNet [63]. We provide details of reward models and evaluation metrics in Tab. 6.
@@ 8.3 Training Details
The loss weight $\lambda$ for reward consistency loss is different for each condition. Specifically, $\lambda$ is 0.5, 0.5, 1.0, 1.0, and 10 for segmentation mask, depth, hed edge, canny edge, and LineArt edge condition, respectively. For all experiments, we first fine-tune the pre-trained ControlNet until convergence using a batch size of 256 and a learning rate of 1e-5. We then employ the same batch size and learning rate for 5k iterations reward fine-tuning. To this end, the valid training samples for reward fine-tuning is $256\times 5,000=1,280,000$. We set threshold $t_{\text{thre}}=200$ of Eq.8 in the main paper for all experiments. Diverging from existing methods that use OpenCV’s [2] implementation of the canny algorithm, we have adopted Kornia’s [42] implementation to make it differentiable. Our codebase is based on the implementation in HuggingFace’s Diffusers [34], and we do not use classifier-free guidance during the reward fine-tuning process following diffusers.
@@ 9 Proof of Equation 7 in the Main Paper
at any timestep $t$ we can use the predicted $\epsilon(x^{\prime}_{t},c_{v},c_{t},t-1)$ to estimate the real noise $\epsilon$ in Eq. 10, and the above equation can be transformed through straightforward algebraic manipulation to the following form:
$x_{t}\approx\sqrt{\bar{\alpha}_{t}}x_{0}+\sqrt{1-\bar{\alpha}_{t}}\epsilon_{\theta}\left(x_{t}^{\prime},c_{v},c_{t},t-1\right),$ (11)
$x_{0}\approx x_{0}^{\prime}=\frac{x_{t}^{\prime}-\sqrt{1-\alpha_{t}}\epsilon_{\theta}\left(x_{t}^{\prime},c_{v},c_{t},t-1\right)}{\sqrt{\alpha_{t}}}.$
To this end, we can obtain the predicted original image $x^{\prime}_{0}$ at any denoising timestep $t$ and use it as the input for reward consistency loss. However, previous work demonstrates that this approximation only yields a smaller error when the time step $t$ is relatively small [18]. Here we find similar results as shown in Fig. 10, which illustrates the predicted $x^{\prime}_{0}$ is significantly different at different timesteps. We kindly encourage readers to refer to Section 4.3 and Figure 5 in the DDPM [18] paper for more experimental results.
10 Conditioning Scale of ControlNet and T2I-Adapter
@@
To simultaneously achieve control based on text prompts and image conditions, existing controllable generation methods perform an addition operation between the image condition features and the text embedding features. The strength of different conditions can be adjusted through a weight value. Hence, an obvious question arises: can better controllability be achieved by increasing the weight of the image condition features? To answer this question, we conduct experiments under different control scales (The weight of image condition feature) in Fig. 11. It shows that naively increasing the control ratio of image conditions does not enhance controllability and may lead to severe image distortion.
@@ 11 Broader Impact and Limitation
In this paper, we leverage visual discriminative models to evaluate and improve the controllability of the text-to-image model. However, there are several interesting directions for further investigation:

### v2

@@ 6 Conclusion
In this paper, we demonstrate from both quantitative and qualitative perspectives that existing works focusing on controllable generation still fail to achieve precise conditional control, leading to inconsistency between generated images and input conditions. To address this issue, we introduce ControlNet++, it explicitly optimizes the consistency between input conditions and generated images using a pre-trained discriminative reward model in a cycle consistency manner, which is different from existing methods that implicitly achieve controllability through latent diffusion denoising. We also propose a novel and efficient reward strategy that calculates consistency loss by adding noise to input images followed by single-step denoising, thus avoiding the significant computational and memory costs associated with sampling from random Gaussian noise. Experimental results under multiple conditional controls show that ControlNet++ significantly improves controllability without compromising image quality and image-text alignment, offering new insights into controllable visual generation.
ControlNet+++++ +: Improving Conditional Controls with Efficient Consistency Feedback
Supplementary Material
Ming Li Taojiannan Yang Huafeng Kuang Jie Wu
[…]
@@ 7 Overview of Supplementary
  • Section 10: More experiments and analysis.
    • Section 10.1: Effectiveness of conditioning scale of existing methods.
    • Section 10.2: Human evaluation on controllability, text guidance and image quaility.
  • Section 11: Discussion of broader impact and limitation.
  • Section 12: More visualization results.
@@ 8.1 Dataset Details
Considering that the training data for ControlNet [63] has not been publicly released, we need to construct our training dataset. In this paper, we adhere to the dataset construction principles of ControlNet [63], which endeavor to select datasets with more accurate conditional conditions wherever possible. Specifically, for the segmentation condition, previous works have provided datasets with accurately labeled segmentation masks [67, 68, 4]. Therefore, we opt to train our model using these accurately labeled datasets following ControlNet [63]. For the Hed, LineArt edge tasks, it is challenging to find datasets with real and accurate annotations. As a result, following ControlNet [63], we train the model using the MultiGen20M dataset [37], which is annotated by models, to address this issue. Regarding the depth task, existing datasets include masks of certain pixels as having unknown depth values, making them incompatible with the current ControlNet pipeline. Therefore, we also adapt the MultiGen20M depth dataset, which is similar to the dataset constructed by ControlNet [63]. In terms of the canny edge task, no human labels are required in the process, so we also adapt the MultiGen20M dataset. We provide details of the datasets in Table 6.
@@
8.2 Reward Model and Evaluation Details
In general, we deliberately choose slightly weaker models as the reward model and opt for stronger models for evaluation. This practice not only ensures the fairness of the evaluation but also helps to determine whether performance improvements result from alignment with the reward model’s preferences or from a genuine enhancement in controllability. While such an approach is feasible for some tasks (Segmentation, Depth), it becomes challenging to implement for others (Hed, Canny, LineArt Edge) due to the difficulty in finding two distinct reward models. In such cases, we use the same model as both the reward model and the evaluation model. We utilize standard evaluation schemes from their respective research fields to evaluate the input conditions and extracted conditions from the generated images, as demonstrated in Section 4.1 of the main paper. We use the same Hed edge detection model and LineArt edge detection model as ControlNet [63]. We provide details of reward models and evaluation in Table 7.
@@ 8.3 Training Details
The loss weight $\lambda$ for reward consistency loss is different for each condition. Specifically, $\lambda$ is 0.5, 0.5, 1.0, 1.0, and 10 for segmentation mask, depth, hed edge, canny edge, and LineArt edge condition, respectively. For all experiments, we first fine-tune the pre-trained ControlNet until convergence using a batch size of 256 and a learning rate of 1e-5. We then employ the same batch size and learning rate for 10k iterations reward fine-tuning. To this end, the valid training samples for reward fine-tuning is $256\times 10,000=2,560,000$. We set threshold $t_{\text{thre}}=200$ of Eq.8 in the main paper for all experiments. Diverging from existing methods that use OpenCV’s [2] implementation of the canny algorithm, we have adopted Kornia’s [42] implementation to make it differentiable. Our codebase is based on the implementation in HuggingFace’s Diffusers [34], and we do not use classifier-free guidance during the reward fine-tuning process following diffusers.
@@ 9 Proof of Equation 7 in the Main Paper
at any timestep $t$ we can use the predicted $\epsilon(x^{\prime}_{t},c_{v},c_{t},t)$ to estimate the real noise $\epsilon$ in Eq. 10, and the above equation can be transformed through straightforward algebraic manipulation to the following form:
$x_{t}\approx\sqrt{\bar{\alpha}_{t}}x_{0}+\sqrt{1-\bar{\alpha}_{t}}\epsilon_{\theta}\left(x_{t}^{\prime},c_{v},c_{t},t\right),$ (11)
$x_{0}\approx x_{0}^{\prime}=\frac{x_{t}^{\prime}-\sqrt{1-\alpha_{t}}\epsilon_{\theta}\left(x_{t}^{\prime},c_{v},c_{t},t\right)}{\sqrt{\alpha_{t}}}.$
To this end, we can obtain the predicted original image $x^{\prime}_{0}$ at any denoising timestep $t$ and use it as the input for reward consistency loss. However, previous work demonstrates that this approximation only yields a smaller error when the time step $t$ is relatively small [18]. Here we find similar results as shown in Figure 10, which illustrates the predicted $x^{\prime}_{0}$ is significantly different at different timesteps. We kindly encourage readers to refer to Section 4.3 and Figure 5 in the DDPM [18] paper for more experimental results.
10 More Experiments
In this section, we provide additional supplements to the experiments discussed in the main paper, including human evaluation on generated data samples on the Segmentation Mask condition in Sec. 10.2, analysis on conditioning scale of existing methods such as ControlNet [63] and T2I-Adapter [30] in Sec. 10.1.
10.1 Effectiveness of Conditioning Scale
@@
To simultaneously achieve control based on text prompts and image conditions, existing controllable generation methods perform an addition operation between the image condition features and the text embedding features. The strength of different conditions can be adjusted through a weight value. Hence, an obvious question arises: can better controllability be achieved by increasing the weight of the image condition features? To answer this question, we conduct experiments under different control scales (The weight of image condition feature) in Figure 11. It demonstrates that naively increasing the control ratio of image conditions does not enhance controllability and may lead to severe image distortion.
10.2 Human Evaluation
Following ControlNet, we use a single condition for human evaluation. We ask 20 users (12 in ControlNet paper) to select the best image based on three distinct criteria as shown in Table 8. Our ControlNet++ offers better controllability without sacrificing image quality or text guidance.
Table 8: Win rate on ADE20K validation dataset (Segmentation).
20 annotators in total Ours ControlNet T2I-Adapter UniControl
Image-Mask Alignment 76.8% 16.7% 2.0% 4.5%
Image Quality 26.1% 25.8 % 23.6% 24.5 %
Image-Text Alignment 25.3% 25.1% 24.9% 24.7%
@@ 11 Broader Impact and Limitation
In this paper, we use visual discriminative models to evaluate and improve the controllability of text-to-image models. However, we also realize that this work is still insufficient and discuss the following issues:
@@
Discussion on the necessity of controllability: Controllability is important since it allows users to modify image conditions to achieve more flexible and accurate generation. Take LineArt Edge as an example: (1) Freely generating in foreground may change the appearance (e.g., adding a beard for women) that we usually do not expect. (2) Freely generating in background will damage some applications (e.g., blur background). (3) Global free generating may destroy the overall artistic effect of the input image, such as lighting, composition, contrast, etc. Furthermore, we show in Fig.5 of the main paper that more controllable diffusion can in return improve the performance of discriminative models. Beyond image generation, the controllable conditional generation can also be combined with ID preserving methods to perform controllable image editing.
12 More Visualization